# Understanding and Improving Length Generalization in Hierarchical Sparse Attention Models

### v1

@@ 5.3 In-depth Analysis of RULER Retrieval
Our experiments empirically validate the role of our proposed architectural components in improving both these aspects: 1. The Bypassing Residual Path is crucial for effective information integration. We observe that the bypassing path (‘Bypass=yes‘) consistently and significantly improves final answer accuracy, even when the correct chunks are among selected chunks. This provides strong evidence for our hypothesis that retrieved information requires dedicated modulation before being integrated into the main residual stream. The bypassing design proves to be a highly effective mechanism for this modulation. 2. The Chunk Encoder and CLS token improve retrieval prominence. Our results show a direct positive correlation between using the encoder and CLS token and the rank of the correct chunk. These models have the same total parameter count; Models with one more encoder layer have one less lower-decoder layer. The fact that this architectural trade-off leads to superior performance on retrieval-intensive tasks demonstrates that the encoder is a more parameter-efficient mechanism for learning robust retrieval representations and adding a CLS token to generate landmark can improve the parameter efficiency of the chunk encoders. Based on these findings, we also significantly improve length generalization of RAMba, detailed in Appendix. C.
@@ 5.4 The Role of Sparsity
DRT 1 8 4096 88.12 76.24 93.07 85.81 79.21 77.23 90.10 82.18 82.02 68.54 88.76 79.77
DRT 1 8 8192 97.50 84.00 87.00 89.50 97.50 75.00 65.00 79.17 95.00 45.00 45.00 61.67
DRT 1 64 4096 78.18 90.91 96.36 88.48 51.31 76.36 94.55 74.07 50.91 69.09 89.09 69.70
The effectiveness of our model hinges on a nuanced application of sparsity. We identify and analyze three critical dimensions of sparsity, using our best-performing DRT configuration for a controlled study (Table 3).
@@
This dimension refers to the sequence length available to the sparse attention mechanism during training. A longer context provides a larger and more diverse pool of chunk candidates, enabling the model to learn a more robust retrieval policy through contrastive learning. Our results confirm its importance: models trained on a 4K context extrapolate successfully up to 8000x their training length, whereas models trained on a 1K context fail to generalize. Increasing the training length to 8K yields further improvements, establishing a clear trend.
@@
This controls how many of the available chunks the model can attend to at each step. A large Top-K (e.g., 64) allows near-dense global access, while a small Top-K (e.g., 8) enforces a sparse retrieval policy, forcing the model to be highly selective. Enforcing this sparsity is crucial for generalization. At extreme lengths (32M), the model trained with Top-K=8 significantly outperforms its Top-K=64 counterpart, as the sparse policy learned during training better matches the demands of inference on extremely long sequences.
@@
This dimension defines how often the retrieval mechanism (TPR) is invoked—every token (per-token) or less frequently (e.g., every 8 tokens). While less frequent retrieval (per-8-tokens) shows marginal gains on the simple single-retrieval task (S-N), its performance degrades substantially on complex tasks requiring multiple lookups (MQ-N, VT). Therefore, per-token retrieval proves to be the more robust and general-purpose choice, which we use in our final model.
@@ 6 Conclusion
To ensure the reproducibility of our findings, we have provided comprehensive details of our experimental setup. The proposed architectural modifications are described thoroughly, with a detailed architectural diagram and conceptual overview in Figure 1. All hyperparameters and training procedures, including the specific recipes for pre-training and supervised fine-tuning on the RULER and BabiLong benchmarks, are thoroughly documented in Appendix B. The datasets used are publicly available, and our evaluation setup is described in Section 5.1. Furthermore, we commit to releasing our source code upon de-anonymization to allow for full verification of our results and to support future research in this area.
@@
Appendix A Additional Related Work
@@
A.1 Static Sparse Attention
@@
Static sparse attention methods reduce computational complexity by restricting each token to a predefined, fixed subset of other tokens. These data-independent patterns are highly efficient but lack the flexibility required for length extrapolation. Early work like Sparse Transformer (Child et al., 2019) introduced fixed strided and dilated patterns, a foundational concept that influenced many subsequent models. This idea evolved into more complex topologies that combine local windowed attention with some form of global connectivity. For instance, Longformer (Beltagy et al., 2020) and ETC (Ainslie et al., 2020) designated a few “global” tokens to act as information hubs for the entire sequence. BigBird (Zaheer et al., 2021) generalized this by combining local, global, and random connections to approximate the small-world properties of a dense graph. These methods, while successfully extending the manageable context length for tasks like long-document processing, are fundamentally tied to the positional relationships learned during training. Their fixed patterns and reliance on learned positional biases prevent them from generalizing to sequence lengths orders of magnitude beyond what they were trained on.
A.2 Dynamic Sparse Attention
@@
Unlike static methods, dynamic sparse attention mechanisms determine attention patterns adaptively based on the input content, aiming to approximate the expressiveness of full attention by focusing computation on relevant tokens. Early approaches often relied on heuristic-based grouping or clustering. Reformer (Kitaev et al., 2020) pioneered the use of locality-sensitive hashing (LSH) to bucket tokens for attention, while the Routing Transformer (Roy et al., 2020) used online k-means clustering. These methods were a significant step towards content-aware sparsity. Another line of work focused on memory, with Transformer-XL (Dai et al., 2019) introducing a recurrent memory and Memorizing Transformers (Wu et al., 2022) using a kNN index for retrieval from an external long-term memory. While these methods are “dynamic”, their selection mechanisms are still parameterized and optimized for the context lengths seen during training. The learned LSH functions, cluster centroids, or memory management policies are not inherently length-agnostic and thus do not extrapolate well. More recent hardware-aware methods like NSA (Yuan et al., 2025) and MoBA (Lu et al., 2025) have achieved significant practical speedups, but their design goal remains computational efficiency within a known context distribution, not training-free generalization to extreme lengths. Our work, in contrast, focuses specifically on learning a retrieval policy that is robust to drastic changes in sequence length, a distinct and complementary objective.
Appendix B Training Recipe
B.1 Pre-training Recipe
@@
B.2 Supervised Fine-Tuning Recipes
@@
Appendix C Hybrid Architectures with SSM
@@
The recent success of hybrid architectures, which synergistically combine different components to balance performance, efficiency, and long-context reasoning, marks a significant trend in sequence modeling. This paradigm has been prominently validated in models built upon the Mamba architecture, such as Jamba (Lieber et al., 2024), Zamba/Zamba2 (Glorioso et al., 2024b; a), Samba (Ren et al., 2025), and others (Team et al., 2025; Soman et al., 2024). The principle extends to hybrids combining attention with linear RNNs or other mechanisms, as seen in Griffin (De et al., 2024), RWKV-X (Hou et al., 2025), RecurrentGemma (Botev et al., 2024), and large-scale MoE models (MiniMax et al., 2025; Team et al., 2024).
@@
Appendix D LLM Usage Statement

### v2

@@ 5.3 In-depth Analysis of RULER Retrieval
Our experiments empirically validate the role of our proposed architectural components in improving both these aspects: 1. The Bypassing Residual Path is crucial for effective information integration. We observe that the bypassing path (‘Bypass=yes‘) consistently and significantly improves final answer accuracy, even when the correct chunks are among selected chunks. This provides strong evidence for our hypothesis that retrieved information requires dedicated modulation before being integrated into the main residual stream. The bypassing design proves to be a highly effective mechanism for this modulation. 2. The Chunk Encoder and CLS token improve retrieval prominence. Our results show a direct positive correlation between using the encoder and CLS token and the rank of the correct chunk. These models have the same total parameter count; Models with one more encoder layer have one less lower-decoder layer. The fact that this architectural trade-off leads to superior performance on retrieval-intensive tasks demonstrates that the encoder is a more parameter-efficient mechanism for learning robust retrieval representations and adding a CLS token to generate landmark can improve the parameter efficiency of the chunk encoders. Based on these findings, we also significantly improve length generalization of RAMba, detailed in Appendix. D.
@@ 5.4 The Role of Sparsity
SWA+HSA 1 8 4096 88.12 76.24 93.07 85.81 79.21 77.23 90.10 82.18 82.02 68.54 88.76 79.77
SWA+HSA 1 8 8192 97.50 84.00 87.00 89.50 97.50 75.00 65.00 79.17 95.00 45.00 45.00 61.67
SWA+HSA 1 64 4096 78.18 90.91 96.36 88.48 51.31 76.36 94.55 74.07 50.91 69.09 89.09 69.70
The effectiveness of our model critically depends on the careful orchestration of sparsity across multiple dimensions. We conduct a systematic ablation study to analyze three key dimensions of sparsity using our best-performing configuration, with detailed results presented in Table 3.
@@
[…]
@@
This parameter controls the number of chunks the model can attend to at each decoding step, directly regulating the sparsity of the retrieval mechanism. While a large Top-K value (e.g., 64) permits nearly dense global attention, a small Top-K (e.g., 8) enforces strict sparsity, compelling the model to learn highly discriminative retrieval. This enforced sparsity during training proves essential for extreme-length generalization. At 32M tokens, the model trained with Top-K=8 substantially outperforms its Top-K=64 counterpart, demonstrating that a sparse retrieval policy learned during training generalizes more effectively to the resource constraints and scale demands of inference at unprecedented sequence lengths.
@@
This dimension defines the retrieval interval: how many tokens are generated before the HSA mechanism refreshes its selected chunks. In other words, it specifies whether chunk selection is updated every single token (per-token retrieval) or once per block of multiple tokens (per-chunk retrieval). Two competing strategies exist: per-token retrieval offers maximal flexibility through fine-grained, dynamic context access at every step, while per-chunk retrieval potentially provides more stable training signals by reducing update frequency. To empirically determine the optimal strategy, we compare per-token retrieval against per-8-token retrieval. The results show a clear tradeoff: per-8-token retrieval achieves marginal gains on simple single-needle tasks (S-N) but suffers severe degradation on complex multi-query tasks (MQ-N, VT) requiring multiple targeted retrievals. This decisively favors per-token retrieval as the more robust choice for our final architecture.
5.5 Scalability and General capabilities
To further validate the scalability and general capabilities of our proposed architecture, we evaluate a scaled-up Mixture-of-Experts (MoE) variant (8B parameters, 1B active). The detailed training recipe and model specifications follow the setup in Hu et al. (2025c). As shown in Appendix A, our architecture achieves competitive performance on general reasoning benchmarks (e.g., GSM8K, MMLU) compared to standard Transformers, confirming that our length-extrapolation modifications do not compromise short-context capabilities.
@@ 6 Conclusion
Acknowledgements
We thank the anonymous reviewers for their insightful comments. This work was supported by the Ant Group Research Intern Program.
@@ Acknowledgements
To ensure the reproducibility of our findings, we have provided comprehensive details of our experimental setup. The proposed architectural modifications are described thoroughly, with a detailed architectural diagram and conceptual overview in Figure 1. All hyperparameters and training procedures, including the specific recipes for pre-training and supervised fine-tuning on the RULER and BabiLong benchmarks, are thoroughly documented in Appendix C. The datasets used are publicly available, and our evaluation setup is described in Section 5.1. Furthermore, we commit to releasing our source code upon de-anonymization to allow for full verification of our results and to support future research in this area.
@@
Appendix A Evaluation on General Capabilities at Scale
A critical concern for specialized long-context architectures is whether the modifications introduced to support extreme length extrapolation—specifically the hierarchical sparse attention and chunk encoding—negatively impact performance on standard short-context tasks.
To empirically verify the scalability of our approach and ensure no performance regression occurs in general domains, we evaluate a scaled-up version of our model, denoted as SWA-HSA-MoE.
Experimental Setup.
We conduct a controlled comparison between our method and a standard Transformer baseline in the large-scale regime:
  • Model Scale: Both models are 8B-parameter Mixture-of-Experts (MoE) architectures with approximately 1B active parameters per token.
  • Training Data: Both models were pre-trained on 8 trillion tokens using identical data mixtures and training recipes.
  • Architecture: The Transformer-MoE baseline uses standard dense attention with MoE Feed-Forward Networks (FFNs). The SWA-HSA-MoE replaces the attention mechanism with our proposed Hierarchical Sparse Attention (HSA) framework (including the non-linear chunk encoder and bypassing residual path) while retaining the identical MoE FFN structure.
We refer readers to Hu et al. (2025c) for comprehensive details regarding the training infrastructure, MoE implementation, and dataset composition used in these experiments.
Results.
We evaluate both models on a diverse set of standard short-context benchmarks covering knowledge, reasoning, math, and coding. As shown in Table 4, SWA-HSA-MoE achieves competitive or superior performance across the majority of tasks:
  1. 1.
    No Regression: The overall average score of SWA-HSA-MoE (57.69) surpasses the standard Transformer-MoE baseline (55.62), confirming that our long-context mechanisms do not compromise core language modeling capabilities.
  2. 2.
    Reasoning & Coding Gains: Notably, our model shows significant improvements in complex reasoning tasks, including GSM8K (+0.61), MATH (+4.02), and coding benchmarks like MBPP+ (+5.56). This suggests that the hierarchical selection capability may offer benefits even in non-retrieval-heavy contexts by efficiently filtering information.
Table 4: General Capabilities Evaluation (8B MoE). Comparison between a standard Transformer-MoE and our proposed SWA-HSA-MoE. Both models have 8B total parameters (1B active) and are trained on 8T tokens. Our method demonstrates robust performance on standard short-text benchmarks, outperforming the baseline on 6 out of 7 tasks.
Benchmark Category Transformer-MoE SWA-HSA-MoE (Ours)
MMLU Knowledge & Reasoning 58.74 57.83
PIQA Knowledge & Reasoning 77.48 78.84
BBH Knowledge & Reasoning 50.34 51.70
GSM8K Math Reasoning 66.41 67.02
MATH Math Reasoning 37.96 41.98
HumanEval+ Coding 48.17 50.61
MBPP+ Coding 50.26 55.82
Average Overall 55.62 57.69
These results serve as strong empirical evidence that the architectural principles proposed in this work—specifically the synergy between the chunk encoder and bypassing residual path—are scalable and universally effective, extending context length significantly without sacrificing the model’s general utility.
Appendix B Additional Related Work
@@
B.1 Static Sparse Attention
@@
Static sparse attention methods reduce computational complexity by restricting each token to a predefined, fixed subset of other tokens. These data-independent patterns are highly efficient but lack the flexibility required for length extrapolation. Early work like Sparse Transformer (Child et al., 2019) introduced fixed strided and dilated patterns, a foundational concept that influenced many subsequent models. This idea evolved into more complex topologies that combine local windowed attention with some form of global connectivity. For instance, Longformer (Beltagy et al., 2020) and ETC (Ainslie et al., 2020) designated a few “global" tokens to act as information hubs for the entire sequence. BigBird (Zaheer et al., 2020) generalized this by combining local, global, and random connections to approximate the small-world properties of a dense graph. These methods, while successfully extending the manageable context length for tasks like long-document processing, are fundamentally tied to the positional relationships learned during training. Their fixed patterns and reliance on learned positional biases prevent them from generalizing to sequence lengths orders of magnitude beyond what they were trained on.
B.2 Dynamic Sparse Attention
@@
Unlike static methods, dynamic sparse attention mechanisms determine attention patterns adaptively based on the input content, aiming to approximate the expressiveness of full attention by focusing computation on relevant tokens. Early approaches often relied on heuristic-based grouping or clustering. Reformer (Kitaev et al., 2020) pioneered the use of locality-sensitive hashing (LSH) to bucket tokens for attention, while the Routing Transformer (Roy et al., 2021) used online k-means clustering. These methods were a significant step towards content-aware sparsity. Another line of work focused on memory, with Transformer-XL (Dai et al., 2019) introducing a recurrent memory and Memorizing Transformers (Wu et al., 2022) using a kNN index for retrieval from an external long-term memory. While these methods are “dynamic", their selection mechanisms are still parameterized and optimized for the context lengths seen during training. The learned LSH functions, cluster centroids, or memory management policies are not inherently length-agnostic and thus do not extrapolate well. More recent hardware-aware methods like NSA (Yuan et al., 2025) and MoBA (Lu et al., 2025) have achieved significant practical speedups, but their design goal remains computational efficiency within a known context distribution, not training-free generalization to extreme lengths. Our work, in contrast, focuses specifically on learning a retrieval policy that is robust to drastic changes in sequence length, a distinct and complementary objective.
Learned Sparse Attention and KV Cache Sharing.
Two concurrent lines of work share surface-level similarities with our approach. SeerAttention (Gao et al., 2025) uses learned block-level gating to approximate dense attention efficiently, and YOCO (Sun et al., 2024) introduces a decoder-decoder architecture with a shared global KV cache. However, both focus on efficiency within existing long-context models rather than training-free length generalization. Our work investigates a distinct goal: identifying the architectural components that enable robust extrapolation from short training contexts to extreme lengths.
Appendix C Training Recipe
C.1 Pre-training Recipe
@@
C.2 Supervised Fine-Tuning Recipes
@@
Appendix D Hybrid Architectures with SSM
@@
The recent success of hybrid architectures, which synergistically combine different components to balance performance, efficiency, and long-context reasoning, marks a significant trend in sequence modeling. This paradigm has been prominently validated in models built upon the Mamba architecture, such as Jamba (Lieber et al., 2024), Zamba/Zamba2 (Glorioso et al., 2024b; a), Samba (Ren et al., 2025), and others (Tencent Hunyuan Team, 2025; Soman et al., 2024; Wang et al., 2024). The principle extends to hybrids combining attention with linear RNNs or other mechanisms, as seen in Griffin (De et al., 2024), RWKV-X (Hou et al., 2025), RecurrentGemma (Botev et al., 2024), and large-scale MoE models (MiniMax, 2025; Jamba Team, 2024).
@@
Appendix E Efficiency Analysis
In this section, we provide a detailed analysis of the computational efficiency of our proposed Hierarchical Sparse Attention (HSA) mechanism, including theoretical complexity bounds and empirical measurements of inference latency.
E.1 Complexity Analysis
Consider a model with HSA operating on a context of length $L$, with chunk size $C$, top-$K$ selection, and hidden dimension $d$. We analyze the complexity of each component:
Chunk Encoder.
The chunk encoder processes each chunk of size $C$ independently. With $L/C$ chunks in total, the complexity is:
$O\left(\frac{L}{C}\cdot(4Cd^{2}+2C^{2}d)\right)=O\left(L\cdot(4d^{2}+2Cd)\right),$ (7)
which scales linearly with context length $L$.
Top-$K$ Chunk Selection.
The top-$K$ selection computes scores between the query and each chunk’s landmark representation. For a single decoding step, the complexity is:
$O\left(\frac{L}{C}\cdot d\right).$ (8)
The cumulative complexity across all $L$ steps is:
$O\left(L\cdot\frac{L}{C}\cdot d\right)=O\left(\frac{L^{2}d}{C}\right).$ (9)
Hierarchical Sparse Attention.
HSA computes attention over the selected $K$ chunks, each of size $C$. For a single step:
$O\left(K\cdot(4Cd^{2}+2C^{2}d)\right).$ (10)
The cumulative complexity for all $L$ steps is:
$O\left(L\cdot K\cdot(4Cd^{2}+2C^{2}d)\right).$ (11)
Discussion.
Only the chunk selection component involves a quadratic term with respect to $L$. However, compared to full attention’s $O(L^{2}d)$ complexity, this term is reduced by a factor of $1/C$. Furthermore, chunk selection is performed only once in the entire model (at a single layer), making its quadratic contribution minimal in practice.
E.2 Inference Efficiency
To empirically validate the efficiency of our proposed architectures, we measure inference latency for generating 100 tokens across varying prompt lengths. All experiments use a batch size of 16, and we report the time cost excluding the initial prefilling phase.
Table 7: Inference Time Comparison. Time cost (in seconds) for generating 100 tokens at different prompt lengths. RAMba (Mamba-2 + HSA) maintains near-constant latency regardless of context length, while full-attention Transformers exhibit substantial degradation.
Model 4K $\downarrow$ 16K $\downarrow$ 64K $\downarrow$
Transformer (Full Attention) 2.26 8.90 32.12
Mamba-2 2.92 2.82 2.84
RAMba (Mamba-2 + HSA) 3.14 3.05 2.76
[…]
Appendix F LLM Usage Statement